# Learn to Teach: Sample-Efficient Privileged Learning for Humanoid Locomotion over Diverse Terrains

### v2

@@ V-C Mitigation of policy divergence
In traditional TSL paradigms, this problem can be partially addressed by DAGGER or asymmetric learning. In L2T, all these techniques are naturally blended together as a single-stage co-training process, providing an efficient and elegant solution to address the imitation gap and policy divergence issue. We show that in Fig. 4, the walking gait of the student trained by DAGGER is significantly different from the teacher policy, while L2T can faithfully imitate the teacher. Notice the shape of the toe pad from the L2T student, which slightly tilts up from the horizontal plane, accurately mimicking the teacher’s toe. In contrast, the DAGGER-trained student agent has a flat-ground walking gait, with the toe pad parallel to the horizontal plane.
V-D Ablation study on the mixture coefficient
We conduct an ablation study on the critical components to further explore key design choices within our algorithm framework. Fig. 5(a) demonstrates the importance of the mixture coefficient $\alpha_{\text{mix}}$ in the training process. We observe that while the algorithm becomes unstable with a large $\alpha_{\text{mix}}$, which is caused by letting a highly suboptimal student agent inject too many samples, an appropriate chosen $\alpha_{\text{mix}}$ can benefit the overall training process. We hypothesize that the discrepancies between the student and teacher promote exploration within the action space, enabling both agents to learn from a broader region around the teacher’s actions.
@@ VI Hardware Experiments
VI-A Locomotion over challenging terrains
We report the results of deploying our policy in various outdoor environments, as illustrated in Fig. 1 and the supplementary video. These experiments were conducted around a university campus, including pedestrian walkways, wooden bridges, grass-covered hills, beach volleyball courts, and gravel paths. Extensive outdoor testing highlights the policy’s capability to maintain a reliable and natural walking gait and withstand diverse environmental conditions, such as walking on wet surfaces and operating with strong wind.
Surprisingly, the policy shows generalization ability to scenarios not included in the training. For example, on grass hills, due to rain, the grass and the soil underneath exhibit a certain level of deformation upon impact, which increases the difficulty of state estimation and thus further increases the observation noise. Moreover, the policy can walk on the beach volleyball court with sandy terrain, as shown in Fig. 1. Despite the robot not being calibrated or trained for such conditions, the policy adapts to the environmental changes without additional training. This demonstrates the robustness and adaptability of our approach. Next, we examine the policy’s performance on terrains with obstacles. The first test involves a crate of gravel shown in Fig. 6(b). The robot consistently performs stepping-in-place actions with a stable motion. Even when the robot occasionally strikes the crate’s edge, it recovers and resumes normal a stepping gait. In the second test, we deploy the robot on a slippery terrain, where we distribute poppy seeds on a whiteboard shown in Fig. 6(c). Surprisingly, our robot does not exhibit any perceptible shaking motions. In contrast, the company controller provided by Agility Robotics fails to walk over. Furthermore, the policy enables omnidirectional control of the robot. Experiments with velocity commands demonstrate the robot’s directional control, which is crucial for practical deployment. Our policy can sustain some deformation and irregularity but may drift and lose balance on such occasions. Readers can find these experiments in the video.
@@ VI-B Perturbation experiments
We evaluate the policy’s response to external perturbations applied during locomotion. Two scenarios are considered: pushing the robot’s center of mass (CoM) from the front, and back using a stick. The policy demonstrates robustness to withstand frontal and rearward pushes while maintaining a stable walking gait. We could not push the robot to failure even when applying a large force. Additionally, we apply a more substantial perturbation by using a harness to pull the robot with an impulsive force (see Fig. 6(a)). The robot adapts dynamically, exhibiting agile adjustments to its gait to compensate for the external perturbations.
@@ VII Conclusion
We introduced L2T-RL, a novel single-stage learning framework that unifies teacher and student training to address sample inefficiency and enhance real-world performance. Our extensive simulation and hardware experiments demonstrate that L2T-RL achieves robust, agile, and precise locomotion while reducing sample complexity by $50\%$ and dramatically saving training time. These results highlight our contributions to redefining teacher-student learning paradigms and paving the way for practical RL-based robotic systems.

### v3

@@ V-B Mitigation of imitation gap
In L2T, imitation learning and sample mixing create a bidirectional feedback loop. The teacher policy benefits from student-generated samples, while the student continuously tracks the evolving teacher. This approach resembles techniques that augment training data [38] in supervised learning works such as computer vision, where randomization injects out-of-distribution data into the training set. In the context of our approach, the out-of-distribution data, with respect to the teacher policy, are from the student.
Figure 5: Our method can mitigate the imitation gap, enabling accurately track teacher policy behavior.
We show that in Fig. 5, the walking gait of the student trained by DAgger is significantly different from the teacher policy, while L2T can faithfully imitate the teacher. Notice the shape of the toe pad from the L2T student, which slightly tilts up from the horizontal plane, accurately mimicking the teacher’s toe. In contrast, the DAgger-trained student agent has a flat-ground walking gait, with the toe pad parallel to the horizontal plane. Details of the walking motions can be found in our video.
Additionally, we compute the $L_{\text{IL}}$ between the teacher and student policies under three settings: L2T, TSL, and L2T without sample mixing ($\alpha_{\text{mix}}=0$) denoted as L2T w/ No Mix as shown in Fig. 4a. This highlights the discrepancy between the two agents across methods and demonstrates that L2T effectively reduces the imitation gap.
Figure 6: (a) Ablation study on the mixture coefficient. (b) CoM velocity versus command using student policy. (c) xyxy trajectory with $x$ direction command using student policy.
V-C Ablation study
We conduct an ablation study on the critical components to further explore key design choices within our algorithm framework. Fig. 6(a) demonstrates the importance of the mixture coefficient $\alpha_{\text{mix}}$ in the training process. We observe that while the algorithm becomes unstable with a large $\alpha_{\text{mix}}$, which is caused by letting a highly suboptimal student agent inject too many samples, an appropriate chosen $\alpha_{\text{mix}}$ can benefit the overall training process. We hypothesize that the discrepancies between the student and teacher promote exploration within the action space, enabling both agents to learn from a broader region around the teacher’s actions. Additionally, we show in Fig. 4b that choosing $L_{\text{IL}}$ slightly outperforms $L_{\text{IL}}+L_{\text{Asym}}$ in our training setup.
@@ VI Hardware Experiments
VI-A Locomotion over real-world uneven terrain
We report the results of deploying our policy in various outdoor environments, as illustrated in Fig. 1 and the supplementary video. These experiments were conducted around a university campus, including walkways, wooden bridges, grass hills, beach volleyball courts, and gravel paths.
Surprisingly, the policy shows generalization ability to scenarios not included in the training. For example, on grass hills, due to rain, the grass and the soil underneath exhibit a certain level of deformation upon impact, which increases the difficulty of state estimation and thus further increases the observation noise. Moreover, the policy can walk on the beach volleyball court with sandy terrain, as shown in Fig. 1. Despite the robot not being calibrated or trained for such conditions, the policy adapts to the environmental changes without additional training. Next, we examine the policy’s performance on terrains with obstacles. The first test involves a crate of gravel shown in Fig. 7(b). The robot consistently performs stepping-in-place actions with a stable motion. Even when the robot occasionally strikes the crate’s edge, it recovers and resumes normal a stepping gait. In the second test, we deploy the robot on a slippery terrain, where we distribute poppy seeds on a whiteboard shown in Fig. 1. Surprisingly, our robot does not exhibit any perceptible shaking motions. In contrast, the company controller provided by Agility Robotics fails to walk over.
Figure 7: (a) we perturb the robot with a harness. (b) we test the step-in-place motion on a crate of gravel. We let the robot walk over (c) high stairs (77cm + 99cm). (d) hill with gravel.
@@ VI-B Perturbation experiments
We evaluate the policy’s response to external perturbations during locomotion. Two scenarios are considered: pushing the robot’s center of mass (CoM) from the front, and back using a stick. The policy demonstrates robustness to withstand frontal and rearward pushes while maintaining a stable walking gait. Additionally, we apply a more substantial perturbation using a harness to pull the robot with an impulsive force (see Fig. 7(a)). The robot adapts dynamically, exhibiting agile adjustments to compensate for the pulling.
@@ VII Conclusion
We introduced L2T-RL, a novel single-stage learning framework that unifies teacher and student training to address sample inefficiency and enhance real-world performance. Our extensive simulation and hardware experiments demonstrate that L2T-RL achieves robust, and agile locomotion while reducing sample complexity by $50\%$ and thus dramatically saving training time. These results highlight our contributions to redefining teacher-student learning paradigms and paving the way for practical RL-based robotic systems.